Shift Summary for eve shift, 8/sep/99

AG (ag@fnal.gov)
Thu, 9 Sep 1999 00:57:29 -0400 (EDT)

Shift Summary for eve shift, 8/sep/99

On Shift: AA, AG

Hours of beam available at all intensities: 51min
Hours of beam available at nominal intensity: 51min
Hours of E832 data (special runs not included): 0h 0min
Hours of special runs (laser scans, muon runs...): 51min

Accounting of lost time:

Scheduled shutdown at 16:00 (which actually happened at 16:21).

========================================================================

Runs ended this shift:

14465 - 3.5M B06 triggers - BEAM_832_REGOUT
14466 - .8M ecal tags - BEAM_KE3HOG2

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================

Accesses during shift:

o PSh and AG went on access to work on laser.

========================================================================

Overview of shift:

Stable beam till 16:21 when the scheduled shutdown starts.
We finished special run with removed regenerator and took a few spills of
ke3 special run. Beam returned ahead of schedule and we started
BEAM_832_KE3HOG2.
~~
========================================================================

Description of Shift:

run____spill______time_______comment

------------------------------------------------------------------------

14465__303_______1530______Shift starts with special tapes loaded and
___________________________special run in progress.

_______330________1607_____Beam is still present. Stop run 14465.

14466__001________1614_____Set ptkick=100MeV and start run 14466
___________________________BEAM_832_KE3HOG2, looking for
___________________________BO2 prescale to reduce dead-time. Eventually
___________________________turned off trigger B01, and B02 prescale is 1.

14466__008________1621_____Beam is gone. Stop run BEAM_832_KE3HOG2 with
___________________________0.78M ECAL tags. We need 4M, the rest is to be
___________________________taken after beam is back.
___________________________Spill signal is gone as well - so DA is not
___________________________running. We could make our own spill signal,
___________________________but it's been decided that it's not worth
___________________________the trouble.

___________________1640____onsplit request tape QKCC19 be mounted on ktev2
___________________________tps3d2.

___________________1800____AA disabled two trouble channels of the NIM
___________________________crate in RR174 from the SDAQ monitoring. The
___________________________+6V channel is dead and the -6V one is unstable.

___________________2330____Spill signal is back! We start COSMICS run.

___________________2340____MCR calls and promises beam in a few minutes.
___________________________They are actually ahead of schedule.
___________________________We stop COSMICS and start BEAM_832_KE3HOG2
___________________________ahead of time.

14472______________2345____And here it is - beam is back and ktevana
___________________________errors are back too...

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift

@ 100MeV ptkick:

o finish BEAM_832_KE3HOG2 (4M more ECAL tags)
o put reg back in (SET)
o BEAM_832_KE3HOG2 again (4-5M ECAL tags)
o BEAM_832_NOTHRESH (400K ECAL tags)

rump NM4 back to 400MeV ptkick

o nominal BEAM_832